OpenStudy (anonymous):

The area of a larger square is 16 times the area of a smaller square. How many times as long is the base of the larger square than the base of the smaller square?

OpenStudy (amistre64):

isnt double a side equal to 4x the area?

OpenStudy (amistre64):

1^2 = 1 2^2 = 4*1 = 4 4^2 = 4*4 = 16 8^2 = 4*16 = 64 .... hmmm

OpenStudy (mathteacher1729):

Yup. Suppose the area of the big square is 16 square units, that means each side is of length four. If the big area is 16 times the smaller area, that meas the smaller area is just 1 square unit and the length of each side is one unit in length. So the length of one side of the bigger square is 4 times the length of the smaller square. :)